OpenStudy (anonymous):

If the quadratic equation ax^2-2bx+c=0 has two equal roots, which of the following is / are true? A. a,b,c form an arithmetic progression. B. a,b,c form a geometric progression. C. Both roots are b/a

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

say, the common root is k

ganeshie8 (ganeshie8):

k+k = ? k*k = ?

ganeshie8 (ganeshie8):

u familiar with the sum of roots, product of roots ?

ganeshie8 (ganeshie8):

sum of roots = -b/a product of roots = c/a

OpenStudy (anonymous):

OMG, i am weak in sum of roots..

ganeshie8 (ganeshie8):

its okay, use those, and find k+k = ?

OpenStudy (anonymous):

why 2 common roots?

ganeshie8 (ganeshie8):

yea, ax^2-2bx+c=0 sum of roots k+k = --2b/a 2k = 2b/a k = b/a -----------------(1)

ganeshie8 (ganeshie8):

product of roots k*k = c/a k^2 = c/a --------------(2)

ganeshie8 (ganeshie8):

a, b, c a, ak, ak^2 ( using (1) and (2) )

ganeshie8 (ganeshie8):

so, a, b, c are in geometric progression with a common ratio of k

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

oh, i understand now. a,b,c cannot form an arithmetic progression?

ganeshie8 (ganeshie8):

they cant, cuz there is no common difference

ganeshie8 (ganeshie8):

there is oly a common ratio, so GP

OpenStudy (caozeyuan):

I would use an easier method. two roots are equation=determinant is 0, that is, b^2-4ac=0,thus, ac=(b/2)^2,which leads to the conclusion that a,b,c form a GP

OpenStudy (caozeyuan):

sorry two roots are equal, not two roots are equation. Typo is frustrating!

OpenStudy (anonymous):

okay, clear. i want to ask is there any numbers can both form an arithmetic progression and GP.?

OpenStudy (caozeyuan):

let me see, if three numbers are a,b,c then b^2=ac and a+c=2b, is that possible, I don't know!

OpenStudy (anonymous):

okay, thanks:)

OpenStudy (caozeyuan):

wait a minute! (2b)^2=4ac so a+c)^2=4ac so a-c)^2=0 so a=c but a has to be greater than c to form a meaningful AP, so NO, the series does not exist!